OpenStudy (anonymous):

Can someone please tell me how to do this and how to find the answer? I will give a medal In one kind of butterfly, brown color is caused by a dominant allele, B, and a white color is caused by a recessive allele, b,. The table shows the observed numbers of phenotypes in a population of 100 butterflies. Phenotype______Number Of Individuals Brown 84 White 16 What is the frequency of the heterozygous genotype in this population of butterflies? A: 0.24 B: 0.36 C: 0.48 D: 0.60 (I know it is not A or D because I got those answers wrong, and after attempting twice at the question and not getting it correct, I need some help with this.)

OpenStudy (anonymous):

I think its b... NOT COMPLETELY SURE. it think it is though.

OpenStudy (anonymous):

This uses the hardy weinburg equation p^2+2pq+q^2=1 where p= number with dominant allele and q=number with recessive

OpenStudy (anonymous):

therefore q^2=16/100, (that term contains only recessive alleles so you can tell it's the number with bb which is 16 out of the 100) therefore q= 0.4 (square root of 16/100) P+q=1 therefore p=0.6 2pq represents the heterozygous proportion of the population. so 2*0.6*0.4= 0.48 I know it's pretty complicated, so you have any more questions on it, I can try help explain it again :)

OpenStudy (anonymous):

Thank you @Lockolotl